OpenStudy (anonymous):

14. A cartoon plane with four engines can accelerate at 8.9 m/s2 when one engine is running. What is the acceleration of the plane if all four engines are running and each produces the same force?

OpenStudy (anonymous):

F=F1+F2+F3+F4=mA1+mA2+mA3+mA4=m(a1+a2+a3+a4) i f all forces are equal then F=4F1=m(4A1) thus F=4(a1)